OpenStudy (anonymous):

Write the equation of the line that passes through (-1, 4) with a slope of -1 in standard form.

OpenStudy (anonymous):

well u have equation y-y1=m(x-x1) y-4=-1(x-(-1)) y-4=-1(x+1) y+x-3=0

OpenStudy (anonymous):

One format for straight-line equations is called the "point-slope" form. For this one, they give you a point (x1, y1) and a slope m, and have you plug it into this formula: y – y1 = m(x – x1) Don't let the subscripts scare you. They are just intended to indicate the point they give you. You have the generic "x" and generic "y" that are always in your equation, and then you have the specific x and y from the point they gave you; the specific x and y are what is subscripted in the formula. Here's how you use the point-slope formula: